# Bipartite Fluctuations and Charge Fractionalization in Quantum Wires

### v3

@@ Abstract
where $\cal Y$} is the same quantity in the absence of interactions for $K=1$ i.e. $\pi^{2}\frac{({\cal F}_{Q(L_{r})}+{\cal F}_{{\tilde{J}}(L_{r})})}{2}-{\cal Y}=\left(\frac{1}{K}(f_{R}^{2}+f_{L}^{2})-1\right)\ln\frac{L_{r}}{\alpha}=2\gamma\ln\frac{L_{r}}{\alpha}$. Here, $2\gamma$ has the physical interpretation of a dimensionless resistance $r=\frac{Re^{2}}{h}$ and $x=L_{r}$ is the size of each path forming the ring. The form of fractional charges and the Luttinger parameter are measurable through the fit of this formula. The power-law form of mesoscopic interferences i.e. of $\langle{\cal T}\rangle$ is then a marker of the entangled nature of the two counter-propagating fractional wavepackets at zero temperature. When the right-moving charge (wavepacket) $f_{R}=\frac{1+K}{2}$ is measured at $x=L_{r}+\epsilon$ in the region $B$ then it is entangled with the left-moving fractional charge (wavepacket) which has moved backward and therefore is yet in the region $A$ inside the ring. The power-law decay of the interferences can be understood as the probability for the left-fractional charge $f_{L}=\frac{1-K}{2}$ to reach $x=L_{r}\sim ut$ at time $t$ i.e. $|\langle{\cal L}_{L}^{f_{L}}(x,t){\cal L}_{L}^{f_{L}}(0,0)\rangle|_{x=L_{r}\rightarrow ut}^{2}\propto\left(\frac{L_{r}}{a}\right)^{-2\gamma}$ [10].
Fractional Charges and Quantum Spin Chains.— Here, we introduce a correspondence with quantum spin chains from the Jordan-Wigner transformation at half-filling i.e. for $\langle n_{i}\rangle=\frac{1}{2}$. For spin polarized fermions with an interaction between nearest neighboring sites, the equivalent spin Hamiltonian is of the form of an XXZ chain
@@
with the length of the wire $L=N\alpha$. We have the correspondences $S_{i}^{z}=c^{\dagger}_{i}c_{i}-\frac{1}{2}$, $S_{j}^{+}S_{j+1}^{-}=c^{\dagger}_{j}c_{j+1}$. The hopping term of fermions on the lattice becomes an $XY$ spin interaction (with equal couplings along $X$ and $Y$) and the interaction $U$ gives rise to an Ising interaction. When $U\ll 2t$, the low-energy theory of this quantum spin model at half-filling is therefore the Luttinger liquid in Eq. (1) [2, 28]. The electron charge density $n_{i}$ corresponds to the spin variable $S_{i}^{z}$, $n_{i}=S_{i}^{z}+\frac{1}{2}$ with $S_{i}^{z}=\pm\frac{1}{2}$ i.e. $\langle n_{i}\rangle=\frac{1}{2}$. The electron current density operator then precisely corresponds to
$j_{i}=t\tilde{j}_{i}=2t(S_{i}^{x}S_{i-1}^{y}-S_{i}^{y}S_{i-1}^{x})=2t{\cal O}_{i,i-1}^{xy-yx};$ (14)
the $\cal O$} operator coupling sites i,i1i,i-1 is the analogue of the $\tilde{j}_{i}=i(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i})$ current density introduced above. In the Letter [27], the authors introduce a representation of the current which can be written in terms of Majorana operators. We develop the correspondence with the ${\cal O}_{i,i-1}^{xy-yx}$ operator in the Supplementary Material [29].
For free fermions i.e. $U=0$, generalizing the derivations in Refs. [21], we find the exact formula from the $XY$ spin chain for bipartite current fluctuations [29]
@@
We remind the formula obtained for the charge fluctuations from the correlations of $S_{i}^{z}$ i.e. ${\cal F}_{Q(x)}=\frac{1}{\pi^{2}}(\ln\frac{x}{\alpha}+\gamma+\ln 2+1)$ [21, 29]. Adding interactions, this is identical to rescale the quantum fields as $\phi(x)\rightarrow\sqrt{K}\phi(x)$ and $\theta(x)\rightarrow\frac{1}{\sqrt{K}}\theta(x)$ in Eq. (1) compared to the free electron model such that we reproduce the leading form of the correlation functions in Eq. (6). From a correspondence with Bethe Ansatz, we have the identification for the Luttinger parameter $K=\frac{\pi}{2\arccos(-\frac{U}{2t})}$ [30].
In Figs. 1 and 2, we present DMRG results of the formula (9) with the quantum spin chain Hamiltonian at half-filling [31, 32]. For Periodic Boundary Conditions (PBC) charge fluctuations for a subsystem of size $x$ reveal a logarithmic behavior for $U/t=0$ up to and beyond $\sim 2$. We find the Luttinger parameter in Eq. (Bipartite Fluctuations and Charge Fractionalization in Quantum Wires) [21]. For Open Boundary Conditions (OBC) the charge fluctuations reveal an additional Friedel oscillating term [21]. We also verify the Luttinger parameter from the logarithmic term in the current fluctuations which is obtained from a precise protocol [29]; current fluctuations show an additional linear term which reveals the Mott transition in Fig. 2 inset. This method leads to the fit of fractional charges when adding the logarithmic terms of ${\cal F}_{\tilde{\cal J}}$ and ${\cal F}_{\cal Q}$ in Fig. 2. We also show a comparison with Bethe Ansatz.
Figure 1: Current fluctuations from DMRG with PBC for a wire with 300 sites, 20 sweeps. Data are in solid colored lines and the dashed lines represent a fit with a logarithmic term plus a linear term modulo a constant [29]. At weak $U$, results agree with the analytical form in Eq. (6) and when $U$ becomes substantial a linear form develops revealing the Mott transition.
@@
Mott physics.— In Ref. [33], we have shown that bipartite charge fluctuations are useful to study quantum phase transitions. We emphasize here that the bipartite current fluctuations are also a powerful and precise tool to detect the vicinity of insulating or Mott phases associated to the decay of the logarithmic term of ${\cal F}_{\tilde{\cal J}}$ in Fig. 2. In the spin chain correspondence, this occurs when $U>2t$ or $K<\frac{1}{2}$ such that the interaction term proportional to $(c^{\dagger}_{R}c_{L})^{2}+h.c.$ in the quantum field theory, i.e. the Sine-Gordon term $\cos(4\phi\sqrt{K})$ coming from $S_{j}^{z}S_{j+1}^{z}$ becomes relevant [28], i.e. the mode $\phi(x)$ becomes fixed to a uniform value associated to this insulator i.e. a classical antiferromagnet in the spin language. The Heisenberg spin chain corresponds to $U=2t$ i.e. $K=\frac{1}{2}$. Fluctuations in the charge sector remain very small for $U>2t$ [29]. The electron Green’s function decreases exponentially in space $\sim\exp-\frac{x}{\xi}$ as a result of the charge density wave formation and the linear term in the bipartite current fluctuations $\frac{x}{\xi}$ i.e. give an access to the correlation length $\xi$ or inversely the energy gap. A fit with inset of Fig. 2 gives the estimate $\frac{\alpha}{\xi}\sim e^{\frac{U}{t}1.4533}$.
Localization effect from the interface and Jackiw-Rebbi topological model.— The Jackiw-Rebbi model [34] is a general model of interfaces with two media showing a potential difference such that it admits a zero-energy bound state in the solution of the quantum field theory [35]. The zero-energy bound state occurs even if the potential is smoothly varying around the interface located at $x=\frac{L}{2}$ i.e. with a potential of the form $\sim\Delta\tanh\left(\frac{x-\frac{L}{2}}{\alpha}\right)$ [34, 36]. Here, we ask two questions: can such a bound state be observable with our tool e.g. bipartite current fluctuations? Can the zero-energy solution coexist with the fractional charges of the Luttinger liquid? We show below that our analysis can lead to positive answers. Our tool can then be useful to locate interfaces as well.
@@
associated to the Hamiltonian $\sum_{\epsilon}\epsilon c^{\dagger}_{\epsilon}c_{\epsilon}$. It allows us to show how the eigenstate equation $H,c^{\dagger}_{\epsilon}]=\epsilon c^{\dagger}_{\epsilon}$ turns into [29]
@@
We derive a proof of such an equation formulating a correspondence towards a quantum field theory of a 1D superconducting wire with s-wave (or d-wave) symmetry or a Luther-Emery model [37, 38, 39]. Eq. (17) admits the zero-energy state [29]
@@
From bosonization, the interface produces a backscattering term $\sim\frac{\Delta}{2\pi\alpha}\tanh\left(\frac{x-\frac{L}{2}}{\alpha}\right)\cos(2\phi)$ in the Hamiltonian density (close to the interface) [2]. It goes to zero precisely when $x=\frac{L}{2}$ and for $x\neq\frac{L}{2}$ it takes a dual form as the p-wave pairing term [40] in the Kitaev p-wave superconducting wire [41], if we flip the role of the charge field $\phi$ and current field $\theta$. From this sense, this bound state can be seen as topological [35], i.e. is stable in the presence of interactions. We emphasize here that the peaks in the density probability and in the current fluctuations remain visible for strong interactions $U\sim t$ and similarly for the logarithmic scaling at small $x$ [29].
@@
Conclusion.— We have derived the relation between fractional charges in 1D quantum wires and bipartite chiral charge fluctuations, asumming bipartite charge and current fluctuations. This also provides a way to measure them through the mesoscopic interferences at zero temperature. We have shown a spin chain correspondence and verified the occurrence of fractional charges through DMRG. We have also shown how bipartite current fluctuations represent a useful marker to localize Mott transitions through a linear term. Introducing a smoothly varying potential at the interface between the two sides of the wire, it is yet possible to detect a localized bound state in the presence of interactions coexisting with the fractional charges. Bipartite charge, spin fluctuations and also bipartite current fluctuations may be measured from correlation functions with a double summation on space. We observe progress in developing tools e.g. in ultra-cold atoms [42]. Similarly, coupling a metallic gate along a region of the wire would also allow to resolve information on charge dynamics similarly as in a quantum dot [43]. Our bipartite fluctuations [21] may also be measured from the dynamical spin correlations similarly as for the measure of the quantum Fisher information [25, 26]. Tools are also developed to measure fractional charges in fractional quantum Hall states [14, 15, 16, 17, 22, 23].

### v4

@@ III Fractional Charges, Bipartite Fluctuations and Dephasing in a Ring
where $\cal Y$} is the same quantity in the absence of interactions for $K=1$ i.e. $\pi^{2}\frac{({\cal F}_{Q(L_{r})}+{\cal F}_{{\tilde{J}}(L_{r})})}{2}-{\cal Y}=\left(\frac{1}{K}(f_{R}^{2}+f_{L}^{2})-1\right)\ln\frac{L_{r}}{\alpha}=2\gamma\ln\frac{L_{r}}{\alpha}$. Here, $2\gamma$ has the physical interpretation of a dimensionless resistance $r=\frac{Re^{2}}{h}$ and $x=L_{r}$ is the size of each path forming the ring. The dimensionless resistance comes from the analogy between the effect of intrinsic interactions and the noise coming from a metallic gate with a resistance $R$ [13, 10]. For ballistic quantum wires, $\frac{h}{e^{2}}$ is the unit of resistance justifying the expression ‘dimensionless resistance’ above. The form of fractional charges and the Luttinger parameter are measurable through the fit of this formula. The power-law form of mesoscopic interferences i.e. of $\langle{\cal T}\rangle$ is then a marker of the entangled nature of the two counter-propagating fractional wavepackets at zero temperature. When the right-moving charge (wavepacket) $f_{R}=\frac{1+K}{2}$ is measured at $x=L_{r}+\epsilon$ (with $\epsilon\rightarrow 0$) in the region $B$ then it is entangled with the left-moving fractional charge (wavepacket) which has moved backward and therefore is yet in the region $A$ inside the ring. The power-law decay of the interferences can be understood as the probability for the left-fractional charge $f_{L}=\frac{1-K}{2}$ to reach $x=L_{r}\sim ut$ at time $t$ i.e. $|\langle{\cal L}_{L}^{f_{L}}(x,t){\cal L}_{L}^{f_{L}}(0,0)\rangle|_{x=L_{r}\rightarrow ut}^{2}\propto\left(\frac{L_{r}}{a}\right)^{-2\gamma}$ in the electron Green’s function [10].
Eq. (21) equivalently provides a direct measure of bipartite fluctuations in quantum wires. We emphasize here that the measure of the dephasing factor as a function of length is accessible in a laboratory [12].
IV Fractional Charges and Quantum Spin Chains
From the Jordan-Wigner transformation at half-filling, i.e. for $\langle n_{i}\rangle=\frac{1}{2}$, we present a correspondence for the electron current density operator in quantum spin chains leading then to the measures of fractional charges through DMRG algorithms. For spin polarized fermions with an interaction between nearest neighboring sites, the spin Hamiltonian is an XXZ chain
@@ IV Fractional Charges and Quantum Spin Chains
with the length of the wire $L=N\alpha$. We have the correspondences $S_{i}^{z}=c^{\dagger}_{i}c_{i}-\frac{1}{2}$, $S_{j}^{+}S_{j+1}^{-}=c^{\dagger}_{j}c_{j+1}$. The hopping term of fermions on the lattice becomes an $XY$ spin interaction (with equal couplings along $X$ and $Y$) and the interaction $U$ gives rise to an Ising interaction. When $U<2t$, the low-energy theory of this quantum spin model at half-filling is therefore the Luttinger liquid in Eq. (1) [2, 30]. The electron charge density $n_{i}$ corresponds to the spin variable $S_{i}^{z}$, $n_{i}=S_{i}^{z}+\frac{1}{2}$ with $S_{i}^{z}=\pm\frac{1}{2}$ i.e. $\langle n_{i}\rangle=\frac{1}{2}$. The electron current density operator then precisely corresponds to
$j_{i}=t\tilde{j}_{i}=2t(S_{i}^{x}S_{i-1}^{y}-S_{i}^{y}S_{i-1}^{x})=2t{\cal O}_{i,i-1}^{xy-yx}.$ (23)
We develop on this correspondence in Appendix A. The $\cal O$} operator coupling sites i,i1i,i-1 is the analogue of the $\tilde{j}_{i}=i(c^{\dagger}_{i}c_{i-1}-c^{\dagger}_{i-1}c_{i})$ current density introduced above. In the Letter [31], the authors introduce a representation of the current which can be written in terms of Majorana operators. We develop the correspondence with the ${\cal O}_{i,i-1}^{xy-yx}$ operator in Appendix A.
For free fermions i.e. $U=0$, generalizing the derivations in Refs. [21], we find the exact formula from the $XY$ spin chain for bipartite current fluctuations
@@
For the sake of clarity, we present the derivation of this formula in Appendix A, where we also remind the formula obtained for the charge fluctuations from the correlations of $S_{i}^{z}$ i.e. ${\cal F}_{Q(x)}=\frac{1}{\pi^{2}}(\ln\frac{x}{\alpha}+\gamma+\ln 2+1)$ [21]. Adding interactions, this is identical to rescale the quantum fields as $\phi(x)\rightarrow\sqrt{K}\tilde{\phi}(x)$ and $\theta(x)\rightarrow\frac{1}{\sqrt{K}}\tilde{\theta}(x)$ in Eq. (1) compared to the free electron model such that we reproduce the leading form of the correlation functions in Eq. (14). From a correspondence with Bethe Ansatz, we have the identification for the Luttinger parameter $K=\frac{\pi}{2\arccos(-\frac{U}{2t})}$ [32].
In Figs. 1 and 2, we present DMRG results of the formula (18) with the quantum spin chain Hamiltonian at half-filling [33, 34]. For Periodic Boundary Conditions (PBC) charge fluctuations for a subsystem of size $x$ reveal a logarithmic behavior for $U/t=0$ up to and beyond $\sim 2$. We find the Luttinger parameter in Eq. (III) [21]. For Open Boundary Conditions (OBC) the charge fluctuations reveal an additional Friedel oscillating term [21]. We also verify the Luttinger parameter from the logarithmic term in the current fluctuations which is obtained from a precise protocol (see Appendix A); current fluctuations show an additional linear term which reveals the Mott transition in Fig. 2 inset. This method leads to the fit of fractional charges when adding the logarithmic terms of ${\cal F}_{\tilde{\cal J}}$ and ${\cal F}_{\cal Q}$ in Fig. 2. We also show a comparison with Bethe Ansatz.
Figure 1: Current fluctuations from DMRG with PBC for a wire with 300 sites, 20 sweeps. Data are in solid colored lines and the dashed lines represent a fit with a logarithmic term plus a linear term modulo a constant (see Appendix A). At weak $U$, results agree with the analytical form in Eq. (14) and when $U$ becomes substantial a linear form develops revealing the Mott transition.
@@
In Ref. [35], we have shown that bipartite charge fluctuations are useful to study quantum phase transitions e.g. associated to Mott physics. We emphasize here that the bipartite current fluctuations are also a powerful and precise tool to detect the transition into a Mott phase associated to the decay of the logarithmic term of ${\cal F}_{\tilde{\cal J}}$ in Fig. 2. This occurs when $U>2t$ or $K<\frac{1}{2}$ such that the interaction term $S_{j}^{z}S_{j+1}^{z}$ in the spin language which is proportional to $(c^{\dagger}_{R}c_{L})^{2}+h.c.$ in the quantum field theory, with the electron operators introduced in Eq. (5), gives rise to the relevant Sine-Gordon term $\cos(4\phi\sqrt{K})$ [30], i.e. the mode $\phi(x)$ becomes fixed to a uniform value associated to this insulator i.e. a classical antiferromagnet in the spin language. The Heisenberg spin chain corresponds to $U=2t$ i.e. $K=\frac{1}{2}$. Fluctuations in the charge sector remain very small for $U>2t$ (see Appendix A). The electron Green’s function decreases exponentially in space $\sim\exp-\frac{x}{\xi}$ as a result of the charge density wave formation and the linear term in the bipartite current fluctuations $\frac{x}{\xi}$ i.e. give an access to the correlation length $\xi$ or inversely the energy gap. A fit with inset of Fig. 2 gives the estimate $\frac{\alpha}{\xi}\sim e^{\frac{U}{t}1.4533}$.
V Localization effect from the interface and Jackiw-Rebbi topological model
The Jackiw-Rebbi model [36] is a general model of interfaces with two media showing a potential difference such that it admits a zero-energy bound state in the solution of the quantum field theory [37]. The zero-energy bound state occurs even if the potential is smoothly varying around the interface located at $x=\frac{L}{2}$ i.e. with a potential of the form $\sim\Delta\tanh\left(\frac{x-\frac{L}{2}}{\alpha}\right)$ [36, 38]. Here, we ask two questions: can such a bound state be observable with our tool e.g. bipartite current fluctuations? Can the zero-energy solution coexist with the fractional charges of the Luttinger liquid? We show below that our analysis can lead to positive answers. Our tool can then be useful to locate interfaces as well.
@@ V Localization effect from the interface and Jackiw-Rebbi topological model
associated to the Hamiltonian $\sum_{\epsilon}\epsilon c^{\dagger}_{\epsilon}c_{\epsilon}$. It allows us to show how the eigenstate equation $H,c^{\dagger}_{\epsilon}]=\epsilon c^{\dagger}_{\epsilon}$ turns into
@@
In Appendix B, we derive a proof of such an equation formulating a correspondence towards a quantum field theory of a 1D superconducting wire with s-wave (or d-wave) symmetry or a Luther-Emery model [39, 40, 41]. Eq. (26) admits the zero-energy state (see Appendix B)
@@
From bosonization, the interface produces a backscattering term (proportional to $c^{\dagger}_{L}c_{R}+h.c.$) i.e. $\sim\frac{\Delta}{2\pi\alpha}\tanh\left(\frac{x-\frac{L}{2}}{\alpha}\right)\cos(2\phi)$ in the Hamiltonian density (close to the interface) [2]. It goes to zero precisely when $x=\frac{L}{2}$ and for $x\neq\frac{L}{2}$ it takes a dual form as the p-wave pairing term [42] in the Kitaev p-wave superconducting wire [43], if we flip the role of the charge field $\phi$ and current field $\theta$. From this sense, this bound state can be seen as topological [37], i.e. is stable in the presence of interactions. We emphasize here that the peaks in the density probability and in the current fluctuations remain visible for strong interactions $U\sim t$ and similarly for the logarithmic scaling at small $x$ (see Appendix B).
@@
VI Conclusion
We have derived the relation between fractional charges in 1D quantum wires and bipartite chiral charge fluctuations, asumming bipartite charge and current fluctuations. This also provides a way to measure them through the mesoscopic interferences at zero temperature. We have shown a spin chain correspondence and verified the occurrence of fractional charges through DMRG. We have also shown how bipartite current fluctuations represent a useful marker to localize Mott transitions through a linear term. Introducing a smoothly varying potential at the interface between the two sides of the wire, it is yet possible to detect a localized bound state in the presence of interactions coexisting with the fractional charges. Bipartite charge, spin fluctuations and also bipartite current fluctuations may be measured from correlation functions with a double summation on space. We observe progress in developing tools e.g. in ultra-cold atoms [44]. Similarly, coupling a metallic gate along a region of the wire would also allow to resolve information on charge dynamics similarly as in a quantum dot [45]. Our bipartite fluctuations [21] may also be measured from the dynamical spin correlations similarly as for the measure of the quantum Fisher information [28, 29]. Tools are also developed to measure fractional charges in fractional quantum Hall states [14, 15, 16, 17, 22, 23].